OpenStudy (freckles):

Does anyone know how to show \[e=2+\frac{1}{1+\frac{1}{2+\frac{2}{3+\frac{3}{4+\frac{4}{5+\cdots}}}}}\]

OpenStudy (freckles):

Like showing \[\sqrt{2}=1+\frac{1}{2+\frac{1}{2+\frac{1}{2+\frac{1}{2+\cdots }}}}\] is easy like I know that I can called that whole continued fraction thingy x but then I see that whole continued fraction thingy again inside so I know I can do this: \[x=1+\frac{1}{1+x} \\ x-1=\frac{1}{x+1} \\ (x-1)(x+1)-1=0 \\ x^2-1-1=0 \\ x^2-2=0 \\ x=\sqrt{2} \text{ since } x>0\] but this one seems harder because the top number numbers are changing each time

OpenStudy (freckles):

a long with the side number

OpenStudy (freckles):

like this one is nothing like the other continued fraction I don't think I can use the same method

ganeshie8 (ganeshie8):

e is not algebraic, so i think the same method of solving a polynomial equation wont work for e

OpenStudy (freckles):

http://webserv.jcu.edu/math//vignettes/continued.htm in this link, it says euler used this technique I wonder exactly which technique they are referring to

OpenStudy (anonymous):

you can probably do it with pade approximants since they are connected to generalized continued fractions

OpenStudy (jtvatsim):

There's a similar identity and derivation given here http://www.math.binghamton.edu/dikran/478/Ch7.pdf ending at PDF page 14. Not exactly the same identity given by the OP, but both are very cool.

OpenStudy (freckles):

\[e^{-x}=\sum_{i=0}^{\infty} \frac{(-x)^i}{i!} \\ e^{-x}=1-x+\frac{x^2}{2}-\frac{x^3}{3!}+\cdots \\ \alpha_1=1 \\ \alpha_2=\frac{1}{x} \\ \alpha_3=\frac{2}{x^2} \\ \alpha_4=\frac{3!}{x^3} \\ \cdots \\ \alpha_n=\frac{(n-1)!}{x^{n-1}} \\ e^{-x}=\frac{1}{1}+\frac{1^2}{\frac{1}{x}-1}+\frac{\frac{1}{x^2}}{\frac{2}{x^2}-\frac{1}{x}}+\frac{\frac{4}{x^4}}{\frac{3!}{x^3}-\frac{2}{x^2}}+\cdots +\frac{\frac{(n-2)!^2}{x^{2(n-2)}}}{\frac{(n-1)!}{x^{n-1}}-\frac{(n-2)!}{x^{n-2}}}+\cdots\] trying to following the arctan(x) example in 7.6 ... and using some of the theorems above though this so far not making sense anyways... still playing with what they have

ganeshie8 (ganeshie8):

I'm reading euler's formula and it seems the super simple way to get that result : \[a_0+a_0a_1+\cdots+a_0a_1\cdots a_n = \dfrac{a_0}{1-\dfrac{a_1}{1+a_1-\dfrac{a_2}{1+a_2-\dfrac{\vdots}{\vdots +\dfrac{a_{n-1}}{1+a_{n-1}-\dfrac{a_n}{1+a_n}}}}}} \] Proving this is easy as induction works pretty good here

OpenStudy (freckles):

but deriving it will be killer probably :(

ganeshie8 (ganeshie8):

Not at all, Base case, \(n=1\) : \[RHS = \dfrac{a_0}{1-\dfrac{a_1}{1+a_1}} = \dfrac{a_0(1+a_1)}{1+a_1-a_1}=a_0+a_0a_1=LHS\color{green}{\checkmark}\]

OpenStudy (freckles):

\[\text{ suppose } \\ a_0+a_0a_1+\cdots+a_0a_1\cdots a_k = \dfrac{a_0}{1-\dfrac{a_1}{1+a_1-\dfrac{a_2}{1+a_2-\dfrac{\vdots}{\vdots +\dfrac{a_{k-1}}{1+a_{k-1}-\dfrac{a_k}{1+a_k}}}}}} \\ \text{ for some integer } k \ge 0\] so we add to both sides the following: \[a_0a_1 \cdots a_k a_{k+1}\] and then we....hmmm...thinking...

ganeshie8 (ganeshie8):

I hope you see why proving by induction is trivial here. Its just algebra manipulation. Maybe lets use the result first : consider taylor series of \(e^x\), \[e^x = 1+1*x/1 + 1*x/1*x/2+1*x/2*x/2+x^2/3 + \cdots\]

ganeshie8 (ganeshie8):

compare that with the left hand side and simply plug the values : \[\begin{align}e^x&=1+1*x/1+1*x/1*x/2+\cdots\\~\\ & = \dfrac{1}{1-\dfrac{x}{1+x-\dfrac{x/2}{1+x/2-\vdots }}} \end{align}\] plugging in \(x=1\), we do get a continued fraction for \(e\) but oops! this is not what we're looking for hmm

OpenStudy (jtvatsim):

All of these articles make appeals to "equivalence transformations." That has got to be the key to working with these...

ganeshie8 (ganeshie8):

*typo consider taylor series of \(e^x\), \[e^x = 1+1*x/1 + 1*x/1*x/2+1*x/1*x/2*x/3 + \cdots\]

OpenStudy (freckles):

\[e=\frac{1}{1-\frac{1}{1+1-\frac{1}{2+1-\frac{2(1)}{3+1-\frac{3(1)}{4+1-\cdots }}}}}\]

OpenStudy (freckles):

and I guess we would have do another transformation to get that one continued fraction representation they had on that other site

ganeshie8 (ganeshie8):

oh wait, how did u get that

OpenStudy (freckles):

well I used the thingy you typed and just cleared the compound mini fraction in each sub-fraction if that make sense

OpenStudy (freckles):

I assumed the thingy you were trying to type was the thing from wikepedia

OpenStudy (freckles):

and then they actually have the clearing of the fractions next

OpenStudy (freckles):

one fraction was multiplied by 2/2 another as multiplied by 3/3 ... so on...

OpenStudy (freckles):

I think I hate continued fractions they are so hard to write

ganeshie8 (ganeshie8):

that is clever!

OpenStudy (freckles):

it wasn't my cleverness it was wikepedia's

ganeshie8 (ganeshie8):

Oh I missed that from wiki, I only read the statement of theorem and got excited lol

OpenStudy (freckles):

yeah it was from that one link you gave me

OpenStudy (freckles):

but I think I need to play with fractions to see how they actually got that formula/theorem thingy

ganeshie8 (ganeshie8):

Induction step is easy, at least for you it will be easy I'm sure...

OpenStudy (freckles):

ok thanks @ganeshie8 I will just be doing some playing of my own I will ask if I have any questions

OpenStudy (freckles):

And I'm sure I can figure it out with enough playing